# COMPASS: Context-Modulated PID Attention Steering System for Hallucination Mitigation

### v1

@@ 1 Introduction
LLMs exhibit strong reasoning capabilities but often produce contextual hallucinations, where outputs conflict with the input context despite relevant evidence being present [8, 16]. These errors typically arise when the model over-relies on its parametric knowledge or generated history rather than the provided prompt.
Beyond mitigation, COMPASS is designed as a scientific probe of how LLMs use context. Each component,the Context Reliance Score, the classifier, and the PID controller, offers a transparent mapping between internal attention signals and model behavior. Rather than treating interpretability as a post-hoc visualization problem, we embed interpretability in the generation loop itself, allowing real-time observation and modulation of evidence use. This framework provides a principled way to study and steer complex model dynamics. Existing mitigation strategies include contrastive or context-aware decoding [10], which reweight token probabilities using an auxiliary distribution, and attention-based diagnostics such as Lookback Lens [2], which train classifiers on attention ratios to detect hallucinations and then guide decoding through candidate re-ranking. More recent approaches, such as DAGCD [3], intervene directly in the attention mechanism, but often rely on multi-pass decoding or pre-specified head selections, introducing latency and limiting flexibility. Our contributions are four fold:
@@
  • Context Reliance Score (CRS): The logit of the attention mass on context keys (last query row), a reformulation of the “lookback ratio” from Lookback Lens [2], used as an online per-head context-sensitivity signal for dynamic head selection.
@@ 2.1 Problem Setting and Notation
We study contextual hallucinations: unsupported or factually incorrect tokens given a supplied context. Consider an auto-regressive Transformer [13] with $L$ layers and $H$ heads per layer. At generation step $t$, the prompt is partitioned as a fixed context $C$ (tokens 1:|C|1{:}|C|) followed by a fixed question $Q$ (tokens |C|+1:|C|+|Q||C|{+}1{:}|C|{+}|Q|); the model has produced t1t{-}1 output tokens thereafter. Let $\mathcal{K}_{t}=\{1,\ldots,|C|{+}|Q|{+}t{-}1\}$ denote key positions, with context keys $\mathcal{K}_{C}=\{1,\ldots,|C|\}$ and non-context keys $\mathcal{K}_{G}=\{|C|{+}1,\ldots,|C|{+}|Q|{+}t{-}1\}$ (which include the question and past outputs). For head $h$ in layer $\ell$, let $A_{t}(\ell,h)\in\mathbb{R}^{|\mathcal{K}_{t}|}$ be the (causal-masked) attention distribution for the last query row at step $t$, and $Z_{t}(\ell,h)\in\mathbb{R}^{|\mathcal{K}_{t}|}$ its pre-softmax logits.
@@ 4.2 Phase 2: Modulating Attention Heads
Models and Baselines. Modulation is evaluated on LLaMA-2 7B and Mistral 7B. Baselines include (i) identical sampling path with mitigation disabled, (ii) Lookback Lens guided re-ranking, (iii) Contrastive Decoding [6], and (iv) random-head scaling (same $\alpha$ but on a random head subset).
@@ 8 Related Work
LLMs frequently generate fluent but factually incorrect content (“hallucination”). Prior strategies include grounding with retrieval [4], structured knowledge graphs, or external consistency classifiers [16]. Reinforcement learning from human feedback [1] further improves reliability [9]. However, most approaches act post hoc, correcting outputs after hallucinations emerge, rather than intervening in the model’s internal reasoning. Transformer heads vary widely in function and importance [14]. Some heads are strongly tied to factual grounding, while others promote hallucinations [2] Work on pruning and masking [15, 5] shows selective head control can shift model behavior, but interventions are static. Our method instead uses dynamic, classifier-informed modulation, adjusting hallucination-prone heads online during decoding. Control theory, particularly PID feedback, has been widely applied in dynamical systems [12] , optimization [7] , and reinforcement learning [4]. We frame hallucination control as a feedback problem: a classifier monitors drift, while a PID loop gates attention heads in real time. Unlike retrieval-based grounding [11][4] [1] [9] , or static head pruning [14, 15], our contribution is the first closed-loop framework for hallucination mitigation that unifies detection and internal modulation via PID control.

### v2

@@ 1 Introduction
LLMs exhibit strong reasoning capabilities but often produce contextual hallucinations, where outputs conflict with the input context despite relevant evidence being present [matarazzo2025, zhao2024]. These errors typically arise when the model over-relies on its parametric knowledge or generated history rather than the provided prompt.
Beyond mitigation, COMPASS is designed as a scientific probe of how LLMs use context. Each component,the Context Reliance Score, the classifier, and the PID controller, offers a transparent mapping between internal attention signals and model behavior. Rather than treating interpretability as a post-hoc visualization problem, we embed interpretability in the generation loop itself, allowing real-time observation and modulation of evidence use. This framework provides a principled way to study and steer complex model dynamics. Existing mitigation strategies include contrastive or context-aware decoding [shi2023], which reweight token probabilities using an auxiliary distribution, and attention-based diagnostics such as Lookback Lens [chuang2024], which train classifiers on attention ratios to detect hallucinations and then guide decoding through candidate re-ranking. More recent approaches, such as DAGCD [huang2025], intervene directly in the attention mechanism, but often rely on multi-pass decoding or pre-specified head selections, introducing latency and limiting flexibility. Our contributions are four fold:
@@
  • Context Reliance Score (CRS): The logit of the attention mass on context keys (last query row), a reformulation of the “lookback ratio” from Lookback Lens [chuang2024], used as an online per-head context-sensitivity signal for dynamic head selection.
@@ 2.1 Problem Setting and Notation
We study contextual hallucinations: unsupported or factually incorrect tokens given a supplied context. Consider an auto-regressive Transformer [vaswani2017] with $L$ layers and $H$ heads per layer. At generation step $t$, the prompt is partitioned as a fixed context $C$ (tokens 1:|C|1{:}|C|) followed by a fixed question $Q$ (tokens |C|+1:|C|+|Q||C|{+}1{:}|C|{+}|Q|); the model has produced t1t{-}1 output tokens thereafter. Let $\mathcal{K}_{t}=\{1,\ldots,|C|{+}|Q|{+}t{-}1\}$ denote key positions, with context keys $\mathcal{K}_{C}=\{1,\ldots,|C|\}$ and non-context keys $\mathcal{K}_{G}=\{|C|{+}1,\ldots,|C|{+}|Q|{+}t{-}1\}$ (which include the question and past outputs). For head $h$ in layer $\ell$, let $A_{t}(\ell,h)\in\mathbb{R}^{|\mathcal{K}_{t}|}$ be the (causal-masked) attention distribution for the last query row at step $t$, and $Z_{t}(\ell,h)\in\mathbb{R}^{|\mathcal{K}_{t}|}$ its pre-softmax logits.
@@ 4.2 Phase 2: Modulating Attention Heads
Models and Baselines. Modulation is evaluated on LLaMA-2 7B and Mistral 7B. Baselines include (i) identical sampling path with mitigation disabled, (ii) Lookback Lens guided re-ranking, (iii) Contrastive Decoding [li2023], and (iv) random-head scaling (same $\alpha$ but on a random head subset).
@@ 8 Related Work
LLMs frequently generate fluent but factually incorrect content (“hallucination”). Prior strategies include grounding with retrieval [lewis2009], structured knowledge graphs, or external consistency classifiers [zhao2024]. Reinforcement learning from human feedback [christiano2023] further improves reliability [ouyang2022]. However, most approaches act post hoc, correcting outputs after hallucinations emerge, rather than intervening in the model’s internal reasoning. Transformer heads vary widely in function and importance [voita2019]. Some heads are strongly tied to factual grounding, while others promote hallucinations [chuang2024] Work on pruning and masking [zhang2022, li2025] shows selective head control can shift model behavior, but interventions are static. Our method instead uses dynamic, classifier-informed modulation, adjusting hallucination-prone heads online during decoding. Control theory, particularly PID feedback, has been widely applied in dynamical systems [astrom2008feedback] , optimization [Lin2021] , and reinforcement learning [lewis2009]. We frame hallucination control as a feedback problem: a classifier monitors drift, while a PID loop gates attention heads in real time. Unlike retrieval-based grounding [shuster2021][lewis2009] [christiano2023] [ouyang2022] , or static head pruning [voita2019, zhang2022], our contribution is the first closed-loop framework for hallucination mitigation that unifies detection and internal modulation via PID control.